40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Aug 25 13:42:40 2020, aaron, Update, Lab Work, optics setup 
    Reply  Wed Aug 26 12:12:22 2020, aaron, Update, Lab Work, optics setup 
       Reply  Fri Aug 28 16:59:59 2020, aaron, Update, Lab Work, optics setup 
          Reply  Mon Aug 31 10:00:25 2020, aaron, Update, Lab Work, optics setup 200828_layout.jpg
Message ID: 2561     Entry time: Tue Aug 25 13:42:40 2020     Reply to this: 2562
Author: aaron 
Type: Update 
Category: Lab Work 
Subject: optics setup 

setting up PSOMA beam path

I took some photos of the existing layout. I'll just take apart the E beam path, and leave the W path unchanged for now as reference.

I moved the E fiber output coupler closer to the edge of the table, to make this path easier to reach.

Hopped around on the laser hysterisis curve for a minute. To optimize the temperature,

  1. Started with TEC on near room temperature, then turned on the laser driver with the current set to maximize power on the meter (S122C)
  2. increased the temperature setpoint (decrease resistance setpoint) slowly until the laser power starts to decrease.
  3. Turn off the laser driver. Decrease temperature setpoint to just below the maximum power setpoint.
  4. Turn on the laser driver. Total power is now 4.6mW, compared to ~2mW with TEC off.
ELOG V3.1.3-